OpenStudy (anonymous):

1) Henry can clean the pool three times as fast as John. Together, they can clean the pool in 2 hours. Which equation can be used to find the number of hours, t, it takes Henry to clean the pool alone? Option A: 1/t+1/2t=3 Option B: 1/t+1/3t=2 Option C: 3/t+3/2t=1 Option D: 2/t+2/3t=1 Any help would be greatly appreciated! c:

OpenStudy (anonymous):

Henry-- 2xs as fast as John John-- x Henry-- 2x x + 2x = 2 j + h =2 hours 2 hours... I'm thinking A, but I am NOT completely sure.... :/

OpenStudy (anonymous):

Let 't' be the time it takes henry to clean the pool Therefore it will take john '2t' The equation is \[\frac{ 1 }{ t }+\frac{ 1 }{ 3t } = \frac{ 1 }{ 2 }\] If you multiply both sides by 2 you should get: \[\frac{ 2 }{ t }+\frac{ 2 }{ 3t }=1\] So that would correspond to answer d.

OpenStudy (anonymous):

correction: '2t' should be '3t'

OpenStudy (anonymous):

Are you there?

OpenStudy (anonymous):

Yes. Muchas gracias! :)

OpenStudy (anonymous):

Sure no problem!